A wave is modeled by the wave function y(x,t) = (0.51m) sin[(2π/7.9m)(x-(14.3m/s)t)]. What is the period of the wave?



Round your answer to 2 decimal places.


Pls do not forget the unit

1
Expert's answer
2022-05-19T05:43:24-0400
"y(x,t) = A \sin[(2π/\lambda)(x-v t)]\\
y(x,t) = (0.51m) \sin[(2π/7.9m)(x-(14.3m/s)t)]"

The period of wave

"T=\lambda/v=(7.9\:\rm m)/(14.3\: m/s)=0.55\:\rm s"